Find the eccentricity of the ellipse (x+2)^2/36+(y-4)^2/11=1
Answer by lwsshak3(11628)

Find the eccentricity of the ellipse (x+2)^2/36+(y-4)^2/11=1
eccentricity=c/a
a^2=36
a=6
b^2=11
c^2=a^2-b^2=36-11=25
c=5
eccentricity=5/6